A ferry leaves a harbor and travels to a resort island at an average speed of 18 mph. On the return trip, because of fog, the ferry travels at an average speed of 12 mph. The total time for the trip is 5 h. How far is the island from the harbor?
Answer by richwmiller(17219)

x+y=5
y=5-x
18x=12(5-x)
18x=60-12x
30x=60
x=2 hours
y=3 hours
18*2=36 miles
12*3=36 miles